# Adversarial Attacks of Vision Tasks in the Past 10 Years: A Survey

### v1

@@ 9. Acknowledgments
- Wang et al. (2021) Xiaosen Wang, Xuanran He, Jingdong Wang, and Kun He. 2021. Admix: Enhancing the transferability of adversarial attacks. In ICCV.
- Wang et al. (2023b) Xunguang Wang, Zhenlan Ji, Pingchuan Ma, Zongjie Li, and Shuai Wang. 2023b. InstructTA: Instruction-Tuned Targeted Attack for Large Vision-Language Models. arXiv preprint arXiv:2312.01886 (2023).
- Wang et al. (2023d) Xinpeng Wang, Xiaoyuan Yi, Han Jiang, Shanlin Zhou, Zhihua Wei, and Xing Xie. 2023d. ToViLaG: Your visual-language generative model is also an evildoer. arXiv preprint arXiv:2312.11523 (2023).
- Wang et al. (2024c) Yu Wang, Xiaogeng Liu, Yu Li, Muhao Chen, and Chaowei Xiao. 2024c. Adashield: Safeguarding multimodal large language models from structure-based attack via adaptive shield prompting. arXiv preprint arXiv:2403.09513 (2024).
- Wang et al. (2004) Zhou Wang, Alan C Bovik, Hamid R Sheikh, and Eero P Simoncelli. 2004. Image quality assessment: from error visibility to structural similarity. TIP (2004).
- Wang et al. (2024b) Zefeng Wang, Zhen Han, Shuo Chen, Fan Xue, Zifeng Ding, Xun Xiao, Volker Tresp, Philip Torr, and Jindong Gu. 2024b. Stop reasoning! when multimodal llms with chain-of-thought reasoning meets adversarial images. arXiv preprint arXiv:2402.14899 (2024).
- Wang et al. (2020b) Zifan Wang, Yilin Yang, Ankit Shrivastava, Varun Rawal, and Zihao Ding. 2020b. Towards frequency-based explanation for robust cnn. arXiv preprint arXiv:2005.03141 (2020).
- Wei et al. (2024a) Alexander Wei, Nika Haghtalab, and Jacob Steinhardt. 2024a. Jailbroken: How does llm safety training fail? NIPS (2024).
- Wei et al. (2024b) Hui Wei, Hao Tang, Xuemei Jia, Zhixiang Wang, Hanxun Yu, Zhubo Li, Shin’ichi Satoh, Luc Van Gool, and Zheng Wang. 2024b. Physical adversarial attack meets computer vision: A decade survey. TPAMI (2024).
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. NIPS (2022).
- Welbl et al. (2021) Johannes Welbl, Amelia Glaese, Jonathan Uesato, Sumanth Dathathri, John Mellor, Lisa Anne Hendricks, Kirsty Anderson, Pushmeet Kohli, Ben Coppin, and Po-Sen Huang. 2021. Challenges in detoxifying language models. arXiv preprint arXiv:2109.07445 (2021).
- Wu et al. (2024) Chen Henry Wu, Jing Yu Koh, Ruslan Salakhutdinov, Daniel Fried, and Aditi Raghunathan. 2024. Adversarial Attacks on Multimodal Agents. arXiv preprint arXiv:2406.12814 (2024).
@@
- Wu et al. (2017) Lei Wu, Zhanxing Zhu, et al. 2017. Towards understanding generalization of deep learning: Perspective of loss landscapes. arXiv preprint arXiv:1706.10239 (2017).
- Wu et al. (2023) Yuanwei Wu, Xiang Li, Yixin Liu, Pan Zhou, and Lichao Sun. 2023. Jailbreaking gpt-4v via self-adversarial attacks with system prompts. arXiv preprint arXiv:2311.09127 (2023).
@@
- Xiao et al. (2018) Chaowei Xiao, Bo Li, Jun-Yan Zhu, Warren He, Mingyan Liu, and Dawn Song. 2018. Generating adversarial examples with adversarial networks. arXiv preprint arXiv:1801.02610 (2018).
@@
- Xie et al. (2017a) Cihang Xie, Jianyu Wang, Zhishuai Zhang, Zhou Ren, and Alan Yuille. 2017a. Mitigating adversarial effects through randomization. arXiv preprint arXiv:1711.01991 (2017).
@@
- Xie et al. (2017b) Cihang Xie, Jianyu Wang, Zhishuai Zhang, Yuyin Zhou, Lingxi Xie, and Alan Yuille. 2017b. Adversarial examples for semantic segmentation and object detection. In ICCV.
@@
- Xie et al. (2019a) Cihang Xie, Yuxin Wu, Laurens van der Maaten, Alan L Yuille, and Kaiming He. 2019a. Feature denoising for improving adversarial robustness. In CVPR.
@@
- Xie et al. (2019b) Cihang Xie, Zhishuai Zhang, Yuyin Zhou, Song Bai, Jianyu Wang, Zhou Ren, and Alan L Yuille. 2019b. Improving transferability of adversarial examples with input diversity. In CVPR.
@@
- Xiong et al. (2022) Yifeng Xiong, Jiadong Lin, Min Zhang, John E Hopcroft, and Kun He. 2022. Stochastic variance reduced ensemble adversarial attack for boosting the adversarial transferability. In CVPR.
@@
- Xu et al. (2023b) Nan Xu, Fei Wang, Ben Zhou, Bang Zheng Li, Chaowei Xiao, and Muhao Chen. 2023b. Cognitive overload: Jailbreaking large language models with overloaded logical thinking. arXiv preprint arXiv:2311.09827 (2023).
@@
- Xu et al. (2023a) Peng Xu, Wenqi Shao, Kaipeng Zhang, Peng Gao, Shuo Liu, Meng Lei, Fanqing Meng, Siyuan Huang, Yu Qiao, and Ping Luo. 2023a. Lvlm-ehub: A comprehensive evaluation benchmark for large vision-language models. arXiv preprint arXiv:2306.09265 (2023).
@@
- Xu (2017) W Xu. 2017. Feature squeezing: Detecting adversarial exa mples in deep neural networks. arXiv preprint arXiv:1704.01155 (2017).
@@
- Yan et al. (2022) Chiu Wai Yan, Tsz-Him Cheung, and Dit-Yan Yeung. 2022. Ila-da: Improving transferability of intermediate level attack with data augmentation. In ICLR.
@@
- Yang et al. (2022) Jinyu Yang, Jiali Duan, Son Tran, Yi Xu, Sampath Chanda, Liqun Chen, Belinda Zeng, Trishul Chilimbi, and Junzhou Huang. 2022. Vision-language pre-training with triple contrastive learning. In CVPR.
@@
- Ye et al. (2024) Qinghao Ye, Haiyang Xu, Jiabo Ye, Ming Yan, Anwen Hu, Haowei Liu, Qi Qian, Ji Zhang, and Fei Huang. 2024. mplug-owl2: Revolutionizing multi-modal large language model with modality collaboration. In CVPR.
@@
- Yin et al. (2019) Dong Yin, Raphael Gontijo Lopes, Jon Shlens, Ekin Dogus Cubuk, and Justin Gilmer. 2019. A fourier perspective on model robustness in computer vision. NIPS (2019).
@@
- Yin et al. (2018) Minghao Yin, Yongbing Zhang, Xiu Li, and Shiqi Wang. 2018. When deep fool meets deep prior: Adversarial attack on super-resolution network. In ACM MM.
@@
- Yin et al. (2024) Zhenfei Yin, Jiong Wang, Jianjian Cao, Zhelun Shi, Dingning Liu, Mukai Li, Xiaoshui Huang, Zhiyong Wang, Lu Sheng, Lei Bai, et al. 2024. Lamm: Language-assisted multi-modal instruction-tuning dataset, framework, and benchmark. NIPS (2024).
- Ying et al. (2024a) Zonghao Ying, Aishan Liu, Xianglong Liu, and Dacheng Tao. 2024a. Unveiling the Safety of GPT-4o: An Empirical Study using Jailbreak Attacks. arXiv preprint arXiv:2406.06302 (2024).
@@
- Ying et al. (2024b) Zonghao Ying, Aishan Liu, Tianyuan Zhang, Zhengmin Yu, Siyuan Liang, Xianglong Liu, and Dacheng Tao. 2024b. Jailbreak Vision Language Models via Bi-Modal Adversarial Prompt. arXiv preprint arXiv:2406.04031 (2024).
@@
- Yu et al. (2015) Fisher Yu, Ari Seff, Yinda Zhang, Shuran Song, Thomas Funkhouser, and Jianxiong Xiao. 2015. Lsun: Construction of a large-scale image dataset using deep learning with humans in the loop. arXiv preprint arXiv:1506.03365 (2015).
@@
- Yu et al. (2023) Jiahao Yu, Xingwei Lin, and Xinyu Xing. 2023. Gptfuzzer: Red teaming large language models with auto-generated jailbreak prompts. arXiv preprint arXiv:2309.10253 (2023).
@@
- Yu et al. (2024) Tianyu Yu, Yuan Yao, Haoye Zhang, Taiwen He, Yifeng Han, Ganqu Cui, Jinyi Hu, Zhiyuan Liu, Hai-Tao Zheng, Maosong Sun, et al. 2024. Rlhf-v: Towards trustworthy mllms via behavior alignment from fine-grained correctional human feedback. In CVPR.
@@
- Yuan et al. (2019) Xiaoyong Yuan, Pan He, Qile Zhu, and Xiaolin Li. 2019. Adversarial examples: Attacks and defenses for deep learning. TNNLS (2019).
- Zagoruyko (2016) Sergey Zagoruyko. 2016. Wide residual networks. arXiv preprint arXiv:1605.07146 (2016).
@@
- Zeng et al. (2021) Yan Zeng, Xinsong Zhang, and Hang Li. 2021. Multi-grained vision language pre-training: Aligning texts with visual concepts. arXiv preprint arXiv:2111.08276 (2021).
- Zhang et al. (2020) Hanwei Zhang, Yannis Avrithis, Teddy Furon, and Laurent Amsaleg. 2020. Walking on the edge: Fast, low-distortion adversarial examples. TIFS (2020).
@@
- Zhang et al. (2024b) Hao Zhang, Wenqi Shao, Hong Liu, Yongqiang Ma, Ping Luo, Yu Qiao, and Kaipeng Zhang. 2024b. Avibench: Towards evaluating the robustness of large vision-language model on adversarial visual-instructions. arXiv preprint arXiv:2403.09346 (2024).
@@
- Zhang et al. (2023c) Jiaming Zhang, Xingjun Ma, Xin Wang, Lingyu Qiu, Jiaqi Wang, Yu-Gang Jiang, and Jitao Sang. 2023c. Adversarial prompt tuning for vision-language models. arXiv preprint arXiv:2311.11261 (2023).
@@
- Zhang et al. (2023a) Renrui Zhang, Jiaming Han, Chris Liu, Peng Gao, Aojun Zhou, Xiangfei Hu, Shilin Yan, Pan Lu, Hongsheng Li, and Yu Qiao. 2023a. Llama-adapter: Efficient fine-tuning of language models with zero-init attention. arXiv preprint arXiv:2303.16199 (2023).
@@
- Zhang et al. (2018) Richard Zhang, Phillip Isola, Alexei A Efros, Eli Shechtman, and Oliver Wang. 2018. The unreasonable effectiveness of deep features as a perceptual metric. In CVPR.
@@
- Zhang et al. (2022) Susan Zhang, Stephen Roller, Naman Goyal, Mikel Artetxe, Moya Chen, Shuohui Chen, Christopher Dewan, Mona Diab, Xian Li, Xi Victoria Lin, et al. 2022. Opt: Open pre-trained transformer language models. arXiv preprint arXiv:2205.01068 (2022).
@@
- Zhang et al. (2023d) Xiaoyu Zhang, Cen Zhang, Tianlin Li, Yihao Huang, Xiaojun Jia, Xiaofei Xie, Yang Liu, and Chao Shen. 2023d. A mutation-based method for multi-modal jailbreaking attack detection. arXiv preprint arXiv:2312.10766 (2023).
@@
- Zhang et al. (2024a) Yongting Zhang, Lu Chen, Guodong Zheng, Yifeng Gao, Rui Zheng, Jinlan Fu, Zhenfei Yin, Senjie Jin, Yu Qiao, Xuanjing Huang, et al. 2024a. SPA-VL: A Comprehensive Safety Preference Alignment Dataset for Vision Language Model. arXiv preprint arXiv:2406.12030 (2024).
@@
- Zhang et al. (2023b) Zhexin Zhang, Leqi Lei, Lindong Wu, Rui Sun, Yongkang Huang, Chong Long, Xiao Liu, Xuanyu Lei, Jie Tang, and Minlie Huang. 2023b. Safetybench: Evaluating the safety of large language models with multiple choice questions. arXiv preprint arXiv:2309.07045 (2023).
@@
- Zhao et al. (2017) Jieyu Zhao, Tianlu Wang, Mark Yatskar, Vicente Ordonez, and Kai-Wei Chang. 2017. Men also like shopping: Reducing gender bias amplification using corpus-level constraints. arXiv preprint arXiv:1707.09457 (2017).
@@
- Zhao et al. (2024) Yunqing Zhao, Tianyu Pang, Chao Du, Xiao Yang, Chongxuan Li, Ngai-Man Man Cheung, and Min Lin. 2024. On evaluating adversarial robustness of large vision-language models. NIPS 36 (2024).
@@
- Zhao et al. (2020) Zhengyu Zhao, Zhuoran Liu, and Martha Larson. 2020. Towards large yet imperceptible adversarial image perturbations with perceptual color distance. In CVPR.
@@
- Zheng et al. (2024) Haonan Zheng, Wen Jiang, Xinyang Deng, and Wenrui Li. 2024. Sample-agnostic Adversarial Perturbation for Vision-Language Pre-training Models. arXiv preprint arXiv:2408.02980 (2024).
@@
- Zheng et al. (2023) Xin Zheng, Yanbo Fan, Baoyuan Wu, Yong Zhang, Jue Wang, and Shirui Pan. 2023. Robust physical-world attacks on face recognition. PR (2023).
@@
- Zhou et al. (2018) Wen Zhou, Xin Hou, Yongjun Chen, Mengyun Tang, Xiangqi Huang, Xiang Gan, and Yong Yang. 2018. Transferable adversarial perturbations. In ECCV.
@@
- Zhou et al. (2023a) Ziqi Zhou, Shengshan Hu, Minghui Li, Hangtao Zhang, Yechao Zhang, and Hai Jin. 2023a. Advclip: Downstream-agnostic adversarial examples in multimodal contrastive learning. In ACM MM.
- Zhou et al. (2023b) Ziqi Zhou, Shengshan Hu, Ruizhi Zhao, Qian Wang, Leo Yu Zhang, Junhui Hou, and Hai Jin. 2023b. Downstream-agnostic adversarial examples. In ICCV.
- Zhu et al. (2023) Deyao Zhu, Jun Chen, Xiaoqian Shen, Xiang Li, and Mohamed Elhoseiny. 2023. Minigpt-4: Enhancing vision-language understanding with advanced large language models. arXiv preprint arXiv:2304.10592 (2023).
- Zong et al. (2024) Yongshuo Zong, Ondrej Bohdal, Tingyang Yu, Yongxin Yang, and Timothy Hospedales. 2024. Safety fine-tuning at (almost) no cost: A baseline for vision large language models. arXiv preprint arXiv:2402.02207 (2024).
- Zou et al. (2023) Andy Zou, Zifan Wang, Nicholas Carlini, Milad Nasr, J Zico Kolter, and Matt Fredrikson. 2023. Universal and transferable adversarial attacks on aligned language models. arXiv preprint arXiv:2307.15043 (2023).

### v2

@@ 9. Acknowledgments
- Wang et al. (2021c) Xiaosen Wang, Xuanran He, Jingdong Wang, and Kun He. 2021c. Admix: Enhancing the transferability of adversarial attacks. In ICCV.
- Wang et al. (2023d) Xinpeng Wang, Xiaoyuan Yi, et al. 2023d. ToViLaG: Your visual-language generative model is also an evildoer. arXiv:2312.11523.
- Wang et al. (2021a) Yajie Wang et al. 2021a. Demiguise attack: Crafting invisible semantic adversarial perturbations with perceptual similarity. arXiv:2107.01396.
- Wang et al. (2024c) Yu Wang et al. 2024c. Adashield: Safeguarding multimodal large language models from structure-based attack via adaptive shield prompting. arXiv:2403.09513.
- Wang et al. (2024d) Z Wang et al. 2024d. Stop reasoning! when multimodal llms with chain-of-thought reasoning meets adversarial images. arXiv:2402.14899.
- Wang et al. (2004) Zhou Wang, Alan C Bovik, Hamid R Sheikh, et al. 2004. Image quality assessment: from error visibility to structural similarity. TIP.
- Wang et al. (2021b) Zhibo Wang, Hengchang Guo, Zhifei Zhang, Wenxin Liu, et al. 2021b. Feature importance-aware transferable adversarial attacks. In ICCV.
- Wang et al. (2020b) Zifan Wang, Yilin Yang, Ankit Shrivastava, et al. 2020b. Towards frequency-based explanation for robust cnn. arXiv:2005.03141.
- Wei et al. (2024a) Alexander Wei, Nika Haghtalab, and Jacob Steinhardt. 2024a. Jailbroken: How does llm safety training fail?. In NIPS.
- Wei et al. (2024b) Hui Wei, Hao Tang, Xuemei Jia, Zhixiang Wang, et al. 2024b. Physical adversarial attack meets computer vision: A decade survey. TPAMI.
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. In NIPS.
- Wei et al. (2023) Shaokui Wei, Mingda Zhang, et al. 2023. Shared adversarial unlearning: Backdoor mitigation by unlearning shared adversarial examples. In NIPS.
- Welbl et al. (2021) Johannes Welbl, Amelia Glaese, Jonathan Uesato, et al. 2021. Challenges in detoxifying language models. arXiv:2109.07445.
- Weng et al. (2020) Cheng-Hsin Weng, Yan-Ting Lee, and Shan-Hung Brandon Wu. 2020. On the trade-off between adversarial and backdoor robustness. In NIPS.
- Wu et al. (2024) Chen Henry Wu, Jing Yu Koh, Ruslan Salakhutdinov, et al. 2024. Adversarial Attacks on Multimodal Agents. arXiv:2406.12814.
@@
- Wu et al. (2017) Lei Wu, Zhanxing Zhu, et al. 2017. Towards understanding generalization of deep learning: Perspective of loss landscapes. arXiv:1706.10239.
@@
- Xiao et al. (2018) Chaowei Xiao, Bo Li, Jun-Yan Zhu, Warren He, et al. 2018. Generating adversarial examples with adversarial networks. arXiv:1801.02610.
@@
- Xie et al. (2017a) Cihang Xie, Jianyu Wang, Zhishuai Zhang, et al. 2017a. Mitigating adversarial effects through randomization. arXiv:1711.01991.
@@
- Xie et al. (2017b) Cihang Xie, Jianyu Wang, Zhishuai Zhang, Yuyin Zhou, et al. 2017b. Adversarial examples for semantic segmentation and object detection. In ICCV.
@@
- Xie et al. (2019a) Cihang Xie, Yuxin Wu, Laurens van der Maaten, Alan L Yuille, et al. 2019a. Feature denoising for improving adversarial robustness. In CVPR.
@@
- Xie et al. (2019b) Cihang Xie, Zhishuai Zhang, Yuyin Zhou, Song Bai, et al. 2019b. Improving transferability of adversarial examples with input diversity. In CVPR.
@@
- Xiong et al. (2022) Yifeng Xiong, Jiadong Lin, et al. 2022. Stochastic variance reduced ensemble adversarial attack for boosting the adversarial transferability. In CVPR.
@@
- Xu et al. (2023b) Nan Xu, Fei Wang, et al. 2023b. Cognitive overload: Jailbreaking large language models with overloaded logical thinking. arXiv:2311.09827.
@@
- Xu et al. (2023a) Peng Xu, Wenqi Shao, et al. 2023a. Lvlm-ehub: A comprehensive evaluation benchmark for large vision-language models. arXiv:2306.09265.
@@
- Xu (2017) W Xu. 2017. Feature squeezing: Detecting adversarial exa mples in deep neural networks. arXiv:1704.01155.
@@
- Yan et al. (2022) Chiu Wai Yan, Tsz-Him Cheung, et al. 2022. Ila-da: Improving transferability of intermediate level attack with data augmentation. In ICLR.
@@
- Yang et al. (2022) Jinyu Yang, Jiali Duan, Son Tran, Yi Xu, Sampath Chanda, et al. 2022. Vision-language pre-training with triple contrastive learning. In CVPR.
@@
- Ye et al. (2024) Qinghao Ye, Haiyang Xu, et al. 2024. mplug-owl2: Revolutionizing multi-modal large language model with modality collaboration. In CVPR.
- Yin et al. (2023) Dong Yin, Raphael Gontijo Lopes, et al. 2023. Jailbreaking gpt-4v via self-adversarial attacks with system prompts. arXiv:2311.09127.
@@
- Yin et al. (2019) Dong Yin, Raphael Gontijo Lopes, Jon Shlens, et al. 2019. A fourier perspective on model robustness in computer vision. In NIPS.
- Yin et al. (2025) Jia-Li Yin, Weijian Wang, Wei Lin, et al. 2025. Adversarial-Inspired Backdoor Defense via Bridging Backdoor and Adversarial Attacks. In AAAI.
@@
- Yin et al. (2018) Minghao Yin, Yongbing Zhang, et al. 2018. When deep fool meets deep prior: Adversarial attack on super-resolution network. In ACM MM.
@@
- Yin et al. (2024) Zhenfei Yin, Jiong Wang, et al. 2024. Lamm: Language-assisted multi-modal instruction-tuning dataset, framework, and benchmark. In NIPS.
@@
- Ying et al. (2024a) Zonghao Ying, Aishan Liu, et al. 2024a. Jailbreak Vision Language Models via Bi-Modal Adversarial Prompt. arXiv:2406.04031.
- Ying et al. (2024b) Zonghao Ying, Aishan Liu, et al. 2024b. Unveiling the Safety of GPT-4o: An Empirical Study using Jailbreak Attacks. arXiv:2406.06302.
@@
- Yu et al. (2015) Fisher Yu et al. 2015. Lsun: Construction of a large-scale image dataset using deep learning with humans in the loop. arXiv:1506.03365.
@@
- Yu et al. (2023) Jiahao Yu, Xingwei Lin, et al. 2023. Gptfuzzer: Red teaming large language models with auto-generated jailbreak prompts. arXiv:2309.10253.
@@
- Yu et al. (2024) Tianyu Yu et al. 2024. Rlhf-v: Towards trustworthy mllms via behavior alignment from fine-grained correctional human feedback. In CVPR.
@@
- Yuan et al. (2019) Xiaoyong Yuan, Pan He, et al. 2019. Adversarial examples: Attacks and defenses for deep learning. TNNLS.
- Zagoruyko (2016) Sergey Zagoruyko. 2016. Wide residual networks. arXiv:1605.07146.
@@
- Zeng et al. (2021) Yan Zeng, Xinsong Zhang, et al. 2021. Multi-grained vision language pre-training: Aligning texts with visual concepts. arXiv:2111.08276.
@@
- Zhang et al. (2024a) Hao Zhang et al. 2024a. Avibench: Towards evaluating the robustness of large vision-language model on adversarial visual-instructions. arXiv:2403.09346.
- Zhang et al. (2020) Hanwei Zhang, Yannis Avrithis, Teddy Furon, et al. 2020. Walking on the edge: Fast, low-distortion adversarial examples. TIFS.
@@
- Zhang et al. (2023c) Jiaming Zhang, Xingjun Ma, Xin Wang, Lingyu Qiu, et al. 2023c. Adversarial prompt tuning for vision-language models. arXiv:2311.11261.
- Zhang et al. (2022b) Jianping Zhang, Weibin Wu, Jen-tse Huang, et al. 2022b. Improving adversarial transferability via neuron attribution-based attacks. In CVPR.
@@
- Zhang et al. (2023b) Renrui Zhang, Jiaming Han, et al. 2023b. Llama-adapter: Efficient fine-tuning of language models with zero-init attention. arXiv:2303.16199.
@@
- Zhang et al. (2018) Richard Zhang, Phillip Isola, Alexei A Efros, et al. 2018. The unreasonable effectiveness of deep features as a perceptual metric. In CVPR.
@@
- Zhang et al. (2022a) Susan Zhang, Stephen Roller, et al. 2022a. Opt: Open pre-trained transformer language models. arXiv:2205.01068.
@@
- Zhang et al. (2023d) Xiaoyu Zhang, Cen Zhang, et al. 2023d. A mutation-based method for multi-modal jailbreaking attack detection. arXiv:2312.10766.
@@
- Zhang et al. (2024b) Yongting Zhang et al. 2024b. SPA-VL: A Comprehensive Safety Preference Alignment Dataset for Vision Language Model. arXiv:2406.12030.
@@
- Zhang et al. (2023a) Zhexin Zhang et al. 2023a. Safetybench: Evaluating the safety of large language models with multiple choice questions. arXiv:2309.07045.
@@
- Zhao et al. (2017) Jieyu Zhao et al. 2017. Men also like shopping: Reducing gender bias amplification using corpus-level constraints. arXiv:1707.09457.
@@
- Zhao et al. (2024) Yunqing Zhao, Tianyu Pang, Chao Du, et al. 2024. On evaluating adversarial robustness of large vision-language models. In NIPS.
@@
- Zhao et al. (2020) Zhengyu Zhao et al. 2020. Towards large yet imperceptible adversarial image perturbations with perceptual color distance. In CVPR.
@@
- Zheng et al. (2024) Haonan Zheng et al. 2024. Sample-agnostic Adversarial Perturbation for Vision-Language Pre-training Models. arXiv:2408.02980.
@@
- Zheng et al. (2023) Xin Zheng, Yanbo Fan, Baoyuan Wu, Yong Zhang, Jue Wang, and Shirui Pan. 2023. Robust physical-world attacks on face recognition. PR.
@@
- Zhou et al. (2018) Wen Zhou, Xin Hou, Yongjun Chen, Mengyun Tang, Xiangqi Huang, et al. 2018. Transferable adversarial perturbations. In ECCV.
@@
- Zhou et al. (2023a) Ziqi Zhou, Shengshan Hu, et al. 2023a. Advclip: Downstream-agnostic adversarial examples in multimodal contrastive learning. In ACM MM.
- Zhou et al. (2023b) Ziqi Zhou, Shengshan Hu, Ruizhi Zhao, Qian Wang, Leo Yu Zhang, et al. 2023b. Downstream-agnostic adversarial examples. In ICCV.
- Zhu et al. (2023) Deyao Zhu et al. 2023. Minigpt-4: Enhancing vision-language understanding with advanced large language models. arXiv:2304.10592.
- Zong et al. (2024) Yongshuo Zong et al. 2024. Safety fine-tuning at (almost) no cost: A baseline for vision large language models. arXiv:2402.02207.
- Zou et al. (2023) Andy Zou, Zifan Wang, et al. 2023. Universal and transferable adversarial attacks on aligned language models. arXiv:2307.15043.